# LeanKAN: A Parameter-Lean Kolmogorov-Arnold Network Layer with Improved Memory Efficiency and Convergence Behavior

### v1

@@ III.1 Parameter-Lean Kolmogorov-Arnold Network (LeanKAN)
We propose LeanKAN, an entirely reformulated KAN layer with addition and multiplication nodes that can resolve these three identified issues in MultKAN. Rather than a standard addition layer $\Phi$ with optional multiplication layers $\Omega$ stacked on top such that $\mathbf{z}=\Omega\mathbf{y}=(\Omega\circ\Phi)\mathbf{x}$ as in MultKAN, we instead incorporate the multiplication layers directly into the computation of the subnode $\mathbf{y}$, replacing certain addition connections with multiplication rather than stacking multiplication on top of them. More specifically, starting from an $n_{l}$-dimensional input $\mathbf{x}$ to the layer $l$ and with an output $\mathbf{z}$ of dimension $n_{l+1}$, we apply the gridded basis functions of Eq. 3 to the input to create the traditional KAN activations as in AddKAN, where the gridding and parameterization are all contained in the $\phi_{l,i,j}$ functions that make up $\Phi_{l}$ (see Eqs. 4 and 5). However, in contrast to Eq. 2 where matrix multiplication immediately sums this $n_{l+1}\times n_{l}$ matrix of activations into the $n_{l+1}$-dimensional output vector, here we split the activation matrix into separate multiplication and addition components. Specifically, we define a single hyperparameter $n^{mu}$ (to differentiate from $n^{m}$ in MultKAN, Eq. 6) dictating the number of multiplication input nodes, with $n^{add}$ defined implicitly via $n_{l}^{add}=n_{l}-n_{l}^{mu}$. Then, the hidden nodes can be expressed by
${y}_{l,i}^{\text{mult}}$ $=\prod_{j=1}^{n_{l}^{mu}}{\phi_{l,i,j}}\left(x_{l,j}\right)~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{for}~{}i\in\{1,2,...,n_{l+1}\}\subset\mathbb{N},$ (7)
${y}_{l,i}^{\text{add}}$ $=\sum_{j=n_{l}^{mu}+1}^{n_{l}}{\phi_{l,i,j}}\left(x_{l,j}\right)~{}~{}~{}~{}~{}\text{for}~{}i\in\{1,2,...,n_{l+1}\}\subset\mathbb{N}.$ (8)
@@
In essence, this formulation splits the inputs into two groups based on the ratio of $n^{mu}$ and $n^{add}$, then independently takes the product and sum within the two groups to eliminate the input dimension $j$ of the activation matrix $\Phi_{l}$ and arrive at the output vector $\mathbf{y}$. We contrast this against MultKAN, where the input dimension $j$ of $\Phi_{l}$ is eliminated via summation only, requiring additional terms in the sublayer dimension (recall MultKAN’s $\mathbf{y}$ sublayer being of higher dimension than the real output $\mathbf{z}$) to facilitate pairwise multiplication. Finally, to arrive at the output $\mathbf{z}$, we simply sum $\mathbf{y}^{\text{mult}}$ and $\mathbf{y}^{\text{add}}$ in the $l^{\text{th}}$ layer,
@@
As previously, $n_{l+1}$ is the number of outputs, making $\mathbf{z}_{l}\in\mathbb{R}^{n_{l+1}}$. A LeanKAN layer with four inputs and three outputs is shown in Fig. 1(C), with the sublayer labeled with $\mathbf{y}^{\text{mult}}$ and $\mathbf{y}^{\text{add}}$ demonstrating an equal split between addition and multiplication subnode operations ($n^{mu}=n^{add}$). To summarize, a MultKAN layer includes a standard AddKAN sublayer stacked with an optional multiplication sublayer. LeanKAN, on the other hand, directly modifies the AddKAN sublayer itself to split the nodes between addition and multiplication.
III.2 Structural Benefits of LeanKAN
@@
This new layer design can be used in the output layer without loss of expressivity. By applying the addition sublayer in Eq. 9 after the combined multiplication sublayer (rather than before, as in MultKAN), we allow all output nodes to experience the effects of both multiplication and addition, regardless of the specific $n^{mu}$ used.
@@
Additionally, while MultKAN inflates the number of parameters included in each multiplication node (relative to a standard addition node) linearly with $k$ in order to perform $k^{\text{th}}$ order multiplication, in LeanKAN we do not perform any augmentation of subnode sizes and directly use the order of multiplication equal to $n^{mu}$. In doing so, we achieve a parameter count for a given node structure (e.g., four inputs and three outputs) that is exactly equal to the parameter count used in a traditional AddKAN, and is strictly smaller than that used in a MultKAN, even with the smallest $k=2$ MultKAN value. See this comparison made explicitly in Fig. 1.
@@
This modification not only preserves a parameter-lean structure, but also eliminates the $k$ hyperparameter, instead rolling it into the single $n^{mu}$ hyperparameter, where the multiplication order is specified directly by the number of input nodes selected for multiplication. We find that this does not penalize the performance of LeanKAN (and in fact appears to enhance its performance), while simplifying the hyperparameter tuning process by eliminating a degree of freedom. Table 1 summarizes these key differendes between MultKAN and LeanKAN.
@@
Finally, we emphasize that unlike most KAN augmentation architectures which implement additional treatments around standard KAN layers such as operator learning [10], physics enforcement [17, 6, 7, 8], or ODE coupling [5], the current methodology is developed as a direct and modular replacement for the standard KAN layers themselves, and can generally be applied to any of these augmented KAN structures as originally done with AddKAN and MultKAN layers. In the section of Experiments and Results, we demonstrate the key benefits of LeanKAN identified here in greater detail. All implementation used in the current work starts from the standard AddKAN repository of [15], with modifications carried out to implement MultKAN and LeanKAN.
IV Experiments and Results
IV.1 LeanKAN as a Multidimensional Function Approximator
@@
In this experiment, we compare the capabilities of MultKAN and LeanKAN when given multidimensional inputs and outputs, and discuss how the streamlined structure of LeanKAN lends it to significantly increased memory efficiency, with sparser parameterizations that can improve training dynamics and interpretability. We design the problem with four input nodes $\mathbf{x}=[x_{1},x_{2},x_{3},x_{4}]^{\intercal}$ and four output nodes $\mathbf{z}=[x_{1}x_{2},x_{3}x_{4},x_{1}x_{2},x_{3}x_{4}]^{\intercal}$. In other words, we try to train KAN approximators for the following model:
@@
where $f$ is the target function to be learned. Here, we construct both MultKAN and LeanKAN with a single layer and a four-point grid using radial basis functions $\psi$ for each activation function $\phi$. The hyperbolic tangent normalization function and swish residual activation functions are implemented [16]. Values for [$x_{1},x_{2},x_{3},x_{4}]^{\intercal}$ are uniformly generated between 0 and 1, with 150 training samples and 50 testing samples selected randomly (and identically across the MultKAN and LeanKAN sample sets). The same ADAM [18] learning rate of $1\cdot 10^{-3}$ is used for both architectures.
IV.1.1 Structural Biases of MultKAN and LeanKAN
@@
$z_{3}$ $=y_{3}\cdot{y_{4}}=\left(\sum_{i=1}^{4}\phi_{3,i}\left(x_{i}\right)\right)\left(\sum_{i=1}^{4}\phi_{4,i}\left(x_{i}\right)\right),$ (13)
$z_{4}$ $=y_{5}\cdot{y_{6}}=\left(\sum_{i=1}^{4}\phi_{5,i}\left(x_{i}\right)\right)\left(\sum_{i=1}^{4}\phi_{6,i}\left(x_{i}\right)\right).$ (14)
Eqs. 1114 show that $z_{1}$ and $z_{2}$ are composed of the addition of univariate functions, while $z_{3}$ and $z_{4}$ take products of two groups of sums (as per $k=2$). Therefore, each node in the output layer possesses different levels of expressivity, and performs different operations on the subnode inputs. While $z_{3}$ and $z_{4}$ can easily reconstruct $x_{1}x_{2}$ and $x_{3}x_{4}$ respectively thanks to the multiplication of activations in Eqs. 13 and 14, $z_{1}$ and $z_{2}$ are effectively treated as AddKAN output nodes with summations only in Eqs. 11 and 12, and thus are not capable of expressing $x_{1}x_{2}$ or $x_{3}x_{4}$. In contrast, LeanKAN (Fig. 2(B)) is designed to have an equal level of expressivity with multiplication and addition for all nodes in the output layer such that
@@
for all $k\in${1,2,3,4}, such that each output performs identical operations on its corresponding set of subnode inputs. Thus, all the nodes in the output layer are equally capable of learning multiplicative operations as well as additive operations. A limitation of this reformulation is that the LeanKAN layer studied here constrains multiplications to be on $x_{1}$ and $x_{2}$ only, with addition only on the remaining $x_{3}$ and $x_{4}$. Therefore, the expression limits are shifted from output in MultKAN (Eq. 1114) to input in LeanKAN (Eq. 15), making only $z_{1}$ and $z_{3}$ learnable with LeanKAN. This tradeoff can be seen visually in the network depictions of Fig. 2(A) and (B), where the nodes that “see” the multiplication operator are highlighted in both layer structures, as well as in the training profiles of Fig. 2(C) and (D), where it is seen that only the multiplication-seeing nodes are able to train properly. In practice, the corresponding limitations of these two networks can be overcome by using multi-layer structures, where the fully-connected nature of each independent layer distributes the multiplication operator across all inputs and outputs. Regardless, we find it important to report these limitations in light of the prevalence of single-layer KAN structures in the literature [1, 2, 5].
@@
Figure 2: Comparison between MultKAN and LeanKAN for single-layer multiplication training example. (A-B) Minimal MultKAN and LeanKAN structures that enable the second order multiplication required in this example ($k=2$ and $n^{a}=2$, or $n^{mu}=2$, respectively). Connections and output node labels with access to the multiplication operator are highlighted in blue or green for MultKAN and LeanKAN, respectively. All activations connected to $z_{3}$ are labeled with either a colored (blue for MultKAN, green for LeanKAN) or grey star for nodes encoding the actual output behavior and dummy nodes, respectively. (C-D) Training dynamics for the same two respective networks, given 3,000 epochs. Note that the legends change between subfigures, corresponding to the change in learnable outputs in (A-B). Strongly overlapping dynamics in (D) lead to only two visible traces.
IV.1.2 Improved Memory Efficiency for Training Performance and Interpretability
@@
In addition to a tradeoff in single-layer expressivity, the newly proposed LeanKAN exhibits significant further benefits (without corresponding limitations) in terms of memory efficiency, accuracy, practitioner-friendly output layer interpretation, and hyperparameter count. As reported in Table 1, memory efficiency is substantially improved with LeanKAN. In a standard AddKAN layer with four inputs and four outputs, each input generates four parameterized activations, which sum into the respective outputs. In MultKAN, as seen in Fig. 2(A), this is increased to six activations per input, as $k=2$ dictates that double the activations are required for the $24-n^{a}=$ 2 multiplication nodes. This change can also be seen in Eqs. 13 and 14, which have double the total activations on the right hand side. These additional activations increase the number of total parameters by $50\%$, from 80 to 120. Notably, this is the most parameter-efficient $k$ hyperparameter possible with MultKAN, with higher values of $k$ linearly increasing this scaling through additional sum terms in the products of Eqs. 13 and 14. In contrast, LeanKAN retains the exact number of activations (and thus parameters) as the corresponding AddKAN: four per input, with 80 total parameters.
@@
The increased parameter count in MultKAN does not add nodes, layers, or other structures that might directly aid in the inference procedure and improve its performance relative to LeanKAN. It instead appears to be spent generating a substantial number of dummy (or non-meaningful) activation function terms. For instance, to express $z_{3}=x_{1}x_{2}$, MultKAN has eight activation functions $\phi$ in Eq. 13 but a majority of them (six out of eight) are not necessary to learn $x_{1}x_{2}$. On the contrary, the LeanKAN architecture only needs four activation functions, half of which contribute to the expression of $x_{1}x_{2}$. Without an explicit need for these bulky distributed multiplication terms (e.g., prior physical knowledge of some unique system), we find their inclusion intuitively unnecessary, especially in the physics modeling applications originally proposed in [2] that typically involve relatively compact functional forms more closely resembling LeanKAN’s sparse Eq. 15 than the sixteen-term products of MultKAN’s Eq. 13-14. Quantitatively, we support this claim by highlighting that even with this reduced number of activation functions and reduced number of parameters, LeanKAN is remarkably able to improve on the performance of MultKAN by over an order of magnitude across all useful metrics (i.e. training and testing losses for the two properly-learned outputs), as seen in Fig. 3 (and identically in Figs. 2(C-D)). Contextualized in the $33\%$ reduction in parameters (80 in LeanKAN vs 120 in MultKAN), this performance improvement is fairly remarkable, and suggests promise for the LeanKAN structure.
@@
The current toy case additionally allows for close inspection of the activation function shapes to provide another angle on this analysis. In Fig. 2(A), the eight MultKAN activations used to reconstruct $z_{3}$ are labeled with a blue or grey star for the meaningful activations and the extraneous “dummy” activations, respectively. The same is done with green and grey stars for the four LeanKAN activations reconstructing $z_{3}$ in Fig. 2(B). In the LeanKAN, the two extraneous activations are effectively zeroed out, leaving the two meaningful activations with clean representations of $-x_{1}$ and $-x_{2}$, which are then multiplied together to arrive neatly at $z_{3}=x_{1}x_{2}$. In contrast, MultKAN’s extraneous activations appear to have found local minima at fixed offsets that have to balance each other out in order to reconstruct $z_{3}$. Specifically, the third activation of $x_{2}$ is, similarly to LeanKAN, a clean identity function. However, this is summed at $y_{3}$ with the third activations of $x_{1}$, $x_{3}$, and $x_{4}$, which unlike in the LeanKAN now have roughly constant offset values that cancel each other out at the $y_{3}$ addition node. Similarly, the fourth activation of $x_{1}$ encodes the useful linear function needed here, but it has a negative offset. To counteract this, the fourth dummy activations of the remaining inputs similarly learn fixed offset values, although this time summing to a slightly positive value to correct the fourth activation of $x_{1}$. The end result here is twofold. Mathematically, LeanKAN is able to rapidly converge its meaningful nodes to the correct behavior, as the dummy nodes are sparse and easily zeroed out; while MultKAN appears to converge more slowly, as the dummy nodes introduce local minima and non-optimal behavior via their bulkier sums and products that hinder the learning of the true underlying model. In terms of qualitative interpretability, LeanKAN’s lean architecture enables more direct interpretation of the activation functions ($z_{3}=x_{1}x_{2}$), while MultKAN’s various offsets and constant activations conceal the true underlying function. While both structures would likely converge well to the trivial output studied here if given enough epochs, LeanKAN is able to do so much faster, with fewer parameters, and in a more interpretable way.
While potentially appearing specific to the case at hand, we believe that the current results hold significant weight in more general and complex learning problems as well. If we return to the case originally proposed in Sec. I, we recall that MultKAN there was suggested to learn $xy$ as $[f_{a}(x)+g_{a}(y)]\cdot[f_{m}(x)+g_{m}(y)]$, finally converging to $f_{a}(x)=x,g_{a}(y)=0,f_{m}(x)=0,$ and $g_{m}(y)=y$. The results of Fig. 2 appear to suggest that the presence of the “dummy” activations $g_{a}(y)=0$ and $f_{m}(x)=0$ impair the expressivity and training dynamics of MultKAN, lending it to slow convergence and counterintuitively nonzero $g_{a}(y)$ and $f_{m}(x)$ functions (i.e. nonzero grey starred activations in Fig. 1(A)) that add computational complexity and visual clutter impairing training and human-readability. In other words, MultKAN’s extraneous bulk means that it does not necessarily snap to the trivial solution even in a trivial training case, which may be of concern for users looking to model more complex phenomena that would likely exacerbate these issues. Its struggle (relative to LeanKAN) to learn the basic multiplication operator in a multidimensional case can be traced directly to specific terms and activations in its formulation, leading us to believe that this relative weakness compared to LeanKAN is not an artifact of a manufactured case, but rather an inherent performance drawback that may scale to other applications. While LeanKAN might not be guaranteed to perform in other applications as well as it did here, we can at least observe that it not only performed better via standard loss metrics, but also appeared to snap rapidly to the intuitive and interpretable result, properly zeroing out the unnecessary grey starred activations in Fig. 1(B).
@@
IV.1.3 Higher Order Multiplication

### v2

@@ 3.1 Parameter-Lean Kolmogorov-Arnold Network (LeanKAN)
We propose LeanKAN, an entirely reformulated KAN layer with addition and multiplication nodes that can resolve these three identified issues in MultKAN. Rather than a standard addition layer $\Phi$ with optional multiplication layers $\Omega$ stacked on top such that $\mathbf{z}=\Omega\mathbf{y}=(\Omega\circ\Phi)\mathbf{x}$ as in MultKAN, we instead incorporate the multiplication layers directly into the computation of the subnode $\mathbf{y}$, where we replace certain addition connections with multiplication. More specifically, starting from an $n_{l}$-dimensional input $\mathbf{x}$ to the layer $l$ and with an output $\mathbf{z}$ of dimension $n_{l+1}$, we apply the gridded basis functions of Eq. 3 to the input to create the traditional KAN activations as in AddKAN, where the gridding and parameterization are all contained in the $\phi_{l,i,j}$ functions that make up $\Phi_{l}$ (see Eqs. 4 and 5). However, in contrast to Eq. 2 where matrix multiplication immediately sums this $n_{l+1}\times n_{l}$ matrix of activations into the $n_{l+1}$-dimensional output vector, here we split the activation matrix into separate multiplication and addition components. We define a single hyperparameter $n^{\text{mu}}$ (to differentiate from $n^{m}$ in MultKAN, Eq. 6) dictating the number of multiplication input nodes, with $n^{\text{add}}$ defined implicitly via $n_{l}^{\text{add}}=n_{l}-n_{l}^{\text{mu}}$. Then, the hidden nodes can be expressed by
${y}_{l,i}^{\text{mult}}$ $=\prod_{j=1}^{n_{l}^{\text{mu}}}{\phi_{l,i,j}}\left(x_{l,j}\right)~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{for}~{}i\in\{1,2,...,n_{l+1}\}\subset\mathbb{N},$ (9)
${y}_{l,i}^{\text{add}}$ $=\sum_{j=n_{l}^{\text{mu}}+1}^{n_{l}}{\phi_{l,i,j}}\left(x_{l,j}\right)~{}~{}~{}~{}~{}\text{for}~{}i\in\{1,2,...,n_{l+1}\}\subset\mathbb{N}.$ (10)
@@
This formulation splits the inputs into two groups based on the ratio of $n^{\text{mu}}$ and $n^{\text{add}}$, then independently takes the product and sum within the two groups to eliminate the input dimension $j$ of the activation matrix $\Phi_{l}$ and arrive at the output vector $\mathbf{y}$. We contrast this against MultKAN, where the input dimension $j$ of $\Phi_{l}$ is eliminated via summation only, requiring additional terms in the sublayer dimension (recall MultKAN’s $\mathbf{y}$ sublayer being of higher dimension than the real output $\mathbf{z}$) to facilitate pairwise multiplication. Finally, to arrive at the output $\mathbf{z}$, we simply sum $\mathbf{y}^{\text{mult}}$ and $\mathbf{y}^{\text{add}}$ in the $l^{\text{th}}$ layer,
@@
As previously, $n_{l+1}$ is the number of outputs, making $\mathbf{z}_{l}\in\mathbb{R}^{n_{l+1}}$. A LeanKAN layer with four inputs and three outputs is shown in Fig. 1(C), with the sublayer labeled with $\mathbf{y}^{\text{mult}}$ and $\mathbf{y}^{\text{add}}$ demonstrating an equal split between addition and multiplication subnode operations ($n^{\text{mu}}=n^{\text{add}}$). To summarize, a MultKAN layer includes a standard AddKAN sublayer stacked with an optional multiplication sublayer. LeanKAN, on the other hand, directly modifies the AddKAN sublayer itself to split the nodes between addition and multiplication.
3.2 Structural Benefits of LeanKAN
@@
This new layer design can be used in the output layer without loss of expressivity. By applying the final addition operation in Eq. 11 after the combined multiplication sublayer, we allow all output nodes to experience the effects of both multiplication and addition, regardless of the specific $n^{\text{mu}}$ used.
@@
Additionally, while MultKAN inflates the number of parameters included in each multiplication node (relative to a standard addition node) linearly with $k$ in order to perform $k^{\text{th}}$ order multiplication, in LeanKAN we do not perform any augmentation of subnode sizes and directly use the order of multiplication equal to $n^{\text{mu}}$. In doing so, we achieve a parameter count for a given node structure (e.g., four inputs and three outputs) that is exactly equal to the parameter count used in a traditional AddKAN, and is strictly smaller than that used in a MultKAN, even with the smallest $k=2$ MultKAN value. See this comparison made explicitly in Fig. 1.
@@
This modification not only preserves a parameter-lean structure, but also eliminates the $k$ hyperparameter, instead rolling it into the single $n^{\text{mu}}$ hyperparameter, where the multiplication order is specified directly by the number of input nodes selected for multiplication. We find that this does not penalize the performance of LeanKAN (and in fact appears to enhance its performance), while simplifying the hyperparameter tuning process by eliminating a degree of freedom. Table 1 summarizes these key differences between MultKAN and LeanKAN.
@@
Finally, we emphasize that in contrast to most KAN augmentation architectures which implement additional treatments around standard KAN layers such as operator learning [1], physics enforcement [28, 18, 7, 9], or ODE coupling [13], the current methodology is developed as a direct and modular replacement for the standard KAN layers themselves, and can generally be applied to any of these augmented KAN structures as originally done with AddKAN and MultKAN layers. In the section of Experiments and Results, we demonstrate the key benefits of LeanKAN identified here in greater detail. All implementation used in the current work starts from the standard Julia language AddKAN repository of [22], with modifications carried out to implement MultKAN and LeanKAN.
4 Experiments and Results
4.1 LeanKAN as a Multidimensional Function Approximator
@@
In this experiment, we compare the capabilities of MultKAN and LeanKAN when given multidimensional inputs and outputs, and discuss how the streamlined structure of LeanKAN lends it to significantly increased memory efficiency, with sparser parameterizations that can improve training dynamics and interpretability. We design the problem with four input nodes $\mathbf{x}=[x_{1},x_{2},x_{3},x_{4}]^{\intercal}$ and four output nodes $\mathbf{z}=[x_{1}x_{2},x_{3}x_{4},x_{1}x_{2},x_{3}x_{4}]^{\intercal}$ and train KAN approximators for the following model:
@@
where $f$ is the target function to be learned. Here, we construct both MultKAN and LeanKAN with a single layer and a four-point grid using radial basis functions $\psi$ for each activation function $\phi$. The hyperbolic tangent normalization function and swish residual activation functions are implemented [25]. Values for [$x_{1},x_{2},x_{3},x_{4}]^{\intercal}$ are uniformly generated between 0 and 1, with 150 training samples and 50 testing samples selected randomly (and identically across the MultKAN and LeanKAN sample sets). The same ADAM [12] learning rate of $1\times 10^{-3}$ is used for both architectures.
4.1.1 Structural Biases of MultKAN and LeanKAN
@@
$z_{3}$ $=y_{3}{y_{4}}=\left(\sum_{i=1}^{4}\phi_{3,i}\left(x_{i}\right)\right)\left(\sum_{i=1}^{4}\phi_{4,i}\left(x_{i}\right)\right),$ (15)
$z_{4}$ $=y_{5}{y_{6}}=\left(\sum_{i=1}^{4}\phi_{5,i}\left(x_{i}\right)\right)\left(\sum_{i=1}^{4}\phi_{6,i}\left(x_{i}\right)\right).$ (16)
Eqs. 1316 show that $z_{1}$ and $z_{2}$ are composed of the addition of univariate functions, while $z_{3}$ and $z_{4}$ take products of two groups of sums (as per $k=2$). Therefore, each node in the output layer possesses different levels of expressivity, and performs different operations on the subnode inputs. While $z_{3}$ and $z_{4}$ can easily reconstruct $x_{1}x_{2}$ and $x_{3}x_{4}$ respectively thanks to the multiplication of activations in Eqs. 15 and 16, $z_{1}$ and $z_{2}$ are effectively treated as AddKAN output nodes with summations only in Eqs. 13 and 14, and thus are not capable of expressing $x_{1}x_{2}$ or $x_{3}x_{4}$. In contrast, LeanKAN (Fig. 2(B)) is designed to have an equal level of expressivity with multiplication and addition for all nodes in the output layer such that
@@
for all $k\in${1,2,3,4}. Thus, all the nodes in the output layer are equally capable of learning multiplicative or additive operations. A limitation of this reformulation is that the LeanKAN layer studied here constrains multiplications to be on $x_{1}$ and $x_{2}$ only, with addition only on the remaining $x_{3}$ and $x_{4}$. Therefore, the expression limits are shifted from output in MultKAN (Eq. 1316) to input in LeanKAN (Eq. 17), making only $z_{1}$ and $z_{3}$ learnable with LeanKAN. This tradeoff can be seen visually in the network depictions of Fig. 2(A) and (B), where the nodes that “see” the multiplication operator are highlighted in both layer structures, as well as in the training profiles of Fig. 2(C) and (D), where it is seen that only the multiplication-seeing nodes are able to train properly. In practice, the corresponding limitations of these two networks can be overcome by using multi-layer structures, where the fully-connected nature of each independent layer distributes the multiplication operator across all inputs and outputs. Regardless, we find it important to report these limitations in light of the prevalence of single-layer KAN structures in the literature [17, 16, 13].
@@
Figure 2: Comparison between MultKAN and LeanKAN for single-layer multiplication training example. (A-B) Minimal MultKAN and LeanKAN structures that enable the second order multiplication required in this example ($k=2$ and $n^{a}=2$, or $n^{\text{mu}}=2$, respectively). Connections and output node labels with access to the multiplication operator are highlighted in blue or green for MultKAN and LeanKAN, respectively. All activations connected to $z_{3}$ are labeled with either a colored (blue for MultKAN, green for LeanKAN) or grey star for nodes encoding the actual output behavior and dummy nodes, respectively. (C-D) Training dynamics for the same two respective networks, given 3,000 epochs. Note that the legends change between subfigures, corresponding to the change in learnable outputs in (A-B). Strongly overlapping dynamics in (D) lead to only two visible traces.
4.1.2 Improved Memory Efficiency for Training Performance and Interpretability
@@
In addition to a tradeoff in single-layer expressivity, the newly proposed LeanKAN exhibits significant further benefits (without corresponding limitations) in terms of memory efficiency, accuracy, practitioner-friendly output layer interpretation, and hyperparameter count. As reported in Table 1, memory efficiency is substantially improved with LeanKAN, thanks to the reduced number of activations in the sublayer. While the MultKAN in this case contains 120 parameters, the LeanKAN achieves the same input-output structure with only 80 (the same number as would be needed for AddKAN). Notably, this is the most parameter-efficient $k$ hyperparameter possible with MultKAN, with higher values of $k$ linearly increasing this scaling through additional sum terms in the products of Eqs. 15 and 16. In contrast, LeanKAN retains the exact number of activations (and thus parameters) as the corresponding AddKAN.
@@
The increased parameter count in MultKAN does not add nodes, layers, or other structures that might otherwise improve performance. It instead appears to be spent generating a substantial number of dummy (or non-meaningful) activation function terms. For instance, to express $z_{3}=x_{1}x_{2}$, MultKAN has eight activation functions $\phi$ in Eq. 15 but a majority of them (six out of eight) are mathematically unnecessary. On the contrary, the LeanKAN architecture only needs four activation functions, half of which contribute to the expression of $x_{1}x_{2}$. Without an explicit need for these bulky distributed multiplication terms (e.g., prior physical knowledge of some unique system), we find their inclusion intuitively unnecessary, especially in the physics modeling applications originally proposed in [16] that typically involve relatively compact functional forms. Quantitatively, we support this claim by highlighting that even with this reduced number of activation functions and reduced number of parameters, LeanKAN is remarkably able to improve on the performance of MultKAN by over an order of magnitude across all useful metrics (i.e. training and testing losses for the two properly-learned outputs), as seen in Fig. 3 (and identically in Figs. 2(C-D)). Contextualized in the $33\%$ reduction in parameters (80 in LeanKAN vs 120 in MultKAN), this performance improvement is notable, and suggests promise for the LeanKAN structure.
@@
The current toy case additionally allows for close inspection of the activation function shapes to provide another angle on this analysis. In Fig. 2(A), the eight MultKAN activations used to reconstruct $z_{3}$ are labeled with a blue or grey star for the meaningful activations and the extraneous “dummy” activations, respectively. The same is done with green and grey stars for the four LeanKAN activations reconstructing $z_{3}$ in Fig. 2(B). In the LeanKAN, the two extraneous activations are effectively zeroed out, leaving the two meaningful activations with clean representations of $-x_{1}$ and $-x_{2}$, which are multiplied together to arrive neatly at $z_{3}=x_{1}x_{2}$. In contrast, MultKAN’s extraneous activations appear to have found local minima at fixed offsets that have to balance each other out in order to reconstruct $z_{3}$. Specifically, the third activation of $x_{2}$ is, similarly to LeanKAN, a clean identity function. However, this is summed at $y_{3}$ with the third activations of $x_{1}$, $x_{3}$, and $x_{4}$, which in contrast to the LeanKAN now have roughly constant offset values that cancel each other out at the $y_{3}$ addition node. Similar behavior is seen in the remaining starred $y_{4}$ inputs. The end result here is twofold. Mathematically, LeanKAN is able to rapidly converge its meaningful nodes to the correct behavior, as the dummy nodes are sparse and easily zeroed out; while MultKAN appears to converge more slowly, as the dummy nodes introduce local minima and non-optimal behavior via their bulkier sums and products that hinder the learning of the true underlying model. In terms of qualitative interpretability, LeanKAN’s lean architecture enables more direct interpretation of the activation functions ($z_{3}=x_{1}x_{2}$), while MultKAN’s various offsets and constant activations conceal the true underlying function. While both structures would likely converge well to the trivial output studied here if given enough epochs, LeanKAN is able to do so much faster, with fewer parameters, and in a more interpretable way.
While potentially appearing specific to the case at hand, we believe that the current results hold significant weight in more general and complex learning problems as well. MultKAN’s struggle (relative to LeanKAN) to learn the basic multiplication operator in a multidimensional case can be traced directly to specific terms and activations in its formulation, implying that this relative weakness is not an artifact of a manufactured case, but rather an inherent performance drawback that may scale to other applications where learned functions contain multiplication operations. We further test the performance of these structures in Sec. 4.2.
@@
4.1.3 Higher Order Multiplication